Given this equation Q - 30pa = 100 and Q + 20px + 100 = 0 calculate equilibrium price, equilibrium quantity demanded and equilibrium quantity supplied


1
Expert's answer
2021-07-02T11:39:50-0400

Solution:

At equilibrium Quantity demanded = Quantity supplied

QD = QS

30pa = price of quantity demanded

20px = price of quantity supplied


Derive quantity demanded equation from inverse demand function given:

Inverse demand function: Q – 30p = 100

Quantity demanded function: QD = 100 + 30P

Derive quantity supplied equation from inverse supply function given:

Inverse supply function: Q + 20p + 100 = 0

Quantity supplied function: QS = -100 – 20P

At equilibrium: QD = QS

100 + 30P = -100 – 20P

100 + 100 = -20P – 30P

200 = -50P

P = -4

Equilibrium price = -4


Substitute in the quantity demanded function to derive the equilibrium quantity demanded:

QD = 100 + 30P

QD = 100 + 30(-4)

QD = 100 – 120

QD = -20


Equilibrium quantity demanded = -20


Substitute in the quantity supplied function to derive the equilibrium quantity supplied:

QS = -100 – 20P

QS = -100 – 20(-4)

QS = -100 + 80

QS = -20


Equilibrium quantity supplied = -20